A Complete Bibliography of Publications in Biometrika for the decade 1920–1929

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

(a + b)^{n} [Pea24d]. (P, χ^{2}) [Pea27c]. 0 [Hal27b]. 1 [Hal27b]. B [Pea24d]. χ^{2} [Pea22b, Pea23c]. η [Pea23d]. ∫_{0}^{θ} cos^{n+1} θ dθ [Wis25a, Wis25b]. N [Chu25, Hal27b, Hal27a, Wis25a, Wis25b]. p [Pea24d].

-Fold [Hal27a]. -Function [Pea24d]. -Methode [And23].

Abraham [Ano25b]. Abruptness [Ano25a, San24a, San24b, San24c]. Account [Wis27]. Actual [Pea28c]. Actuary [Ano28i]. Acute [Eld22]. Addenda [Ano23a]. Addendum [Ano23b]. Adolescence [SSK27]. Adult [Ush20]. Adults [CGN23]. Aetiological [Eld22]. Age [Gra28, Pea23a, Pea24e, Gum25]. Agility [Mus26a]. Albinism [PU29]. Albino [Ush20]. Alexander [Ano26b]. Almena [Eld21]. Alter [Gum25]. Alveolar [Pea25e]. Among [Tho23]. Analysis [Stu27, Toc28, And26, And27]. Ancestry [Til23]. Andvord [Eld22].

Anglosaxon [Mor26b]. Anomaly [PS23]. Anthropologie [Ano25c]. Anthropometric [SSK27]. Anwendung [And23, And26, And27]. Any [Chu26, Mil26, Irw27, Pea24e]. Appendix [PE28]. Application [Pea25g, Pea27a, Wis25a, And26, And27]. Applied [Mor26a]. applies [Ano29c]. Approximate [Wis27]. Approximation [Pea24a]. Archaeology [Sto27]. Area [Mil26, Pea23a]. Arrays [R.23]. Artificial [Sun28]. Artificially [Pea21d]. Ascertaining [Pea24g, Pea28c, Woo29]. Association [Toc28, Woo29]. August [Pea25g]. Australian [Mor27]. Avuncular [Pea26b].

Back [Ano28b, Ano28a, Ano28c, Ano28d, Ano29a]. Badarian [Sto27]. Barton [Pea21a]. Based [Mor27]. Basque [Mor29]. Baye [Pea24c, Pea25a]. Bayes [Bur24, Wis27]. be [Eld22, Rho24]. Beardsley [Eld21]. Bearing [E.25, Pea28d]. been [Pea24g, Rho24]. Being [Hal27b, Mor26a, Pea25g, Toc28]. Bement [Eld21]. Bernoulli [Pea25d]. Between [Hal27b, Har26, P.23, SSK27, Art24]. Bi [Gra28]. Bi-Cristal [Gra28]. Binomial [Cam24, Fri25, Pea24d]. Biometric [CGN23, Mor26a, PD24, Wic25]. Biometrika [Ano24a, Ano25b, Ano25c, Ano26c, Ano26b, Ano28i, Ano29c, Irw29]. Biometry [PT28]. Biserial [New25]. Bivariate [Nar23, Pea25b, Pea28b]. Body [San24d]. Bone [Ste29]. Book [E.25, Eld21, Eld22, I.23, Mor26a, P.23, R.23]. Born [Ore20]. Both [Mer28, Pea28a]. Bounding [Mil26]. Boys [Eld29]. Breeding [War22]. British [HM26, Mor24, Sto27]. Brothers [Pea25g]. Browne [P.24, Til23]. Bruce [Pea24j]. Buchanan [Pea26d]. Burmese [Til21].

C [Pea23a]. Calculation [Ano25a, San24a, San24b, San24c]. Calculus [Ano29c]. can [Rho24]. Cancer [Sto24, Sto25]. Capacity [CGN23, HP29, Hol24, MY23, PS27, Til27]. Carbohydrate [Rab27]. Case [Hal27a, Mus26b, Pea21a, PS23, Wic26]. Cathedral [Rei26]. Cephalic [Ore20, PT24]. Cereals [Ano24c]. Certain [Art24, Hol24, IR27, Mor24, Mus26b, NP28a, NP28b, Pea24f, Rho23, Wis27]. Characteristics [Tap27]. Characters [Art24, Hol24]. Childhood [Eld22]. Children [P.23]. Chinese [Ste29]. Chronology [PT28]. Class [Pea23b]. Class-Index [Pea23b]. closed [Mor21]. Coefficient [Pea21c, Pea24a, Pea26c, PJE29]. Coefficients [Ano25a, Hal27a, New25, PE28, Pea28a, PP22, San24a, San24b, San24c, Smi22, Stu21, Tap27, Tho23]. Components [Mer28]. Concerning [Ano23a, Ano24a]. Conditions [Nar23, P.23]. Congenital [Ush24]. Connected [Pea27a]. Consideration [Wis25b]. Consistent [Cam25a]. Constant [Pea25b]. Constants [Ano20a, PD24, PA28, PA29, Sho29]. Contingency [Kon29]. Continuous [Pea24f]. Contribution [Mor29, Pea28b]. Contributions [Ano27c, Pea25c, SN25]. Corpuscle [Wic25, Wic26]. Corpuscles [Wic26]. Correct [Ano25a, San24c]. Correction [Pea23d]. Corrections [Pea28a].

Correlated [Dod27]. Correlation [Ano27d, Hal27a, Mus26a, Mus26b, New25, Ney26b, Ore20, PE28, Pea20b, Pea21c, PP22, Pea22a, Pea23d, Pea23b, Pea24a, Pea25f, Pea25g, Pea27b, Rho23, Rho25, Smi22, Stu21, Tap27, Tho23, Woo29, And26, And27]. Correlations [Yas25]. Corrigenda [Ano27a]. Cows [Toc28]. Craig [Pea29b]. Crania [Hoo26, Mor24, Sto27]. Cranial [PS27, Til27]. Craniology [HM26, Mor25, Mor26a, Mor26b]. Craniometry [Mor29]. Cristal [Gra28]. Criteria [Dod27, NP28a, NP28b]. Criterion [Irw25b]. Critical [P.23]. Critique [Mor26a]. Crude [Ano25a, San24c]. Curve [Ano21b, Ano25b, Mil26, Pea24b]. Curves [Rom24, Wis25a, Wis26, Wis27].

Dairy [Toc28]. Darnley [Pea28d]. Darstellung [Gum24]. Data [Toc28]. Davis [Eld21]. Dawley [Eld21]. Decimal [LP25]. Definition [Pea25e]. Deformed [Pea21d]. Degree [Har26, Pea25g]. Delicate [Ano29c]. Delinquents [Eld21]. Density [San24d]. Derived [New25]. Determination [Pea25g, Stu21, Til27, Wis25a]. Determined [Pea21c]. Determining [Ano27d, Pea21b]. Development [SSK27]. Deviation [Rid29]. Deviations [Chu25, Chu26, Pea27c, Smi22, Yas25]. Dextrality [WP27, Woo28]. Diabetes [Rab27]. Diameter [Gra28]. Diameters [HP29]. Difference [And23, And26, And27, Irw25a, PE23, Pea24g, Pea25g]. Differences [Ano28l, Ste29]. Differential [Pea27a]. Differenzenmethode [And26, And27]. Digital [PS23]. Digitalis [War24]. Digits [Sun28]. Discoverer [Ano25b]. Discussion [Sop28]. Disease [Eld22]. Dislocation [Ush24]. Distribution [Chu25, Hal27b, Irw27, Mer28, Pea26a, Pea27c, PA28, PA29, PJE29, Rid29, Wis28b]. Distributions [Mus26b, Nar23, Pea28a, Pea24h, Sto25, Tch21, Yas25, Yas26]. Does [Rab27]. Dogs [PU29]. Double [Pea24f]. Doublet [Pea24g]. Dr [Ano28i, Pea23a, Pea29b, Stu21]. Dr. [Ano24a, Wis28a]. Drawn [Chu25, Hal27b, Ney26b, Pea24g, PJE29, Rho24, Sop28, SN25].

each [Ano21b]. Early [Mor26b]. Eastern [Bux20]. Edinburgh [Mor26a]. Editor [Ano24a]. Edmund [Ano26c]. Egypt [Sto27]. Egyptian [Mor25, Ore20]. Elements [E.25]. Ellipsoidal [Wic26]. Eminence [HP29]. England [Mor26b]. English [Hoo26]. Entombed [Rei26]. Equally [Hal27b]. Equation [Pea22a]. Equations [Pea27a, And26, And27]. Errata [Ano27a]. Erratum [Ano25a, San24a, San24b]. Error [Kon29, Pea23b, Pea24h, Pea24a, Stu21, Yas26]. Errors [Ano20a, New25, Pea24b, PM27, Stu27]. Études [Sim29]. Evolution [PD21b]. Examination [Sun28, Ush20]. Examined [Pea25a]. Excavated [Sto27]. Excess [GN25]. Expansions [Hen22, IR27]. expectancy [Gum25]. Expectation [Pea24e, Tch21]. Experimental [New25, Pea25a, Stu21]. Experiments [Sun28]. External [PS27]. Extract [Ano24a]. Extreme [Tip25]. Eye [WP27, Woo28]. Eyeball [Ush20]. Eyes [Ush20].

F [Eld22]. Facial [Sto23]. Factor [PM27]. Factorising [Dod27]. Family [Pea21a]. Farringdon [Hoo26]. Fernald [Eld21]. Fertility [May23]. Fifteen [Pea25b]. Final [Sun28]. Finding [Pea25f]. Finger [Eld20]. Finger-Print [Eld20]. Finite [Irw27, SN25]. First [GN25, LP25, Mor23, Mor26b, PE28, Pea24d, Pea25f, PJE29, Til21]. Fit [Pea22b, Pea23c, Pea27c]. Foetal [Ush20]. Fold [Hal27a]. Follow [Mer28]. Followed [Mor21]. Forms [Nar23]. Formula [Bur24]. Formulae [Fri25, PS27, Ste29]. Formulas [I.23]. Four [Sto23]. Fourfold [New25]. Foxglove [War22]. Francis [Irw25a]. Fraternal [Smi22]. Frequency [Ano20a, Ano21b, Irw27, Mer28, Mus26b, Nar23, Pea28a, Pea24f, Pea24h, Pea25b, Pea28b, PA28, PA29, Rom24, Sho29, Tch21, Wis25a, Wis26, Yas25, Yas26, R.23]. Front [Ano26a, Ano27b, Ano28f, Ano28e, Ano28g, Ano28h, Ano29b]. Function [Pea24d]. Functions [Hen22, IR27, LP25, Sho29, Yas25]. Fundamental [Ano21a, Pea20a]. Further [Ano27c, Irw25a, Ney26a, Pea23c, Pea25c, Pea26a, PS27, Pea27b, Sto25, Wis25b].

Galton [Irw25a, Pea25g]. Généalogie [Sim29]. General [Nar23, Pea21b]. Generalisation [Rom24]. Generalised [PM27, Wis26, Wis28b]. Generations [Sto23]. George [Pea26d]. German [And23, And26, And27, Gum24, Gum25]. Giovanni [Pea28b]. Girls [Eld29]. Given [Pea28c, Rho24]. Gland [SSK27]. Glasgow [Eld29]. Gleichungen [And26, And27]. Goitre [Sto24, Sto25, SSK27]. Goodness [Pea22b, Pea23c, Pea27c]. Gordon [Mor26a]. Greater [Pea24e]. Grip [Art24]. Groupings [Ano27d]. Grundbegriffe [Ano26b]. Grundprobleme [Ano26b]. Gumbel [Ano24a].

H [Eld21]. Halley [Ano26c]. Hand [WP27, Woo28]. Harrower [Mor26a]. Having [Irw27]. Hayes [Eld21]. Head [Har26, Mus26a, Til27]. Health [E.25]. Healthy [CGN23]. Height [Eld29]. Heights [GTW25]. Hemisphere [HP29]. Henry [Pea28d]. Hierarchical [Tho23]. Histological [Ush20]. Historic [HM26, Mor26b]. Historical [P.27, Pea24b]. History [Mor24, Pea20b]. Hokien [Mor26a]. Home [P.23]. Hospitals [GTW25]. Human [May23, PD24, San24d, Ush20]. Hybrid [War24]. Hypergeometric [Cam25b]. Hypergeometrical [Pea24f, Pea24i, Rom25].

II [HP29, Irw27, NP28b, PD21b]. III [Cra29]. Illustration [Ano25b, Ano25c, Ano26c, Ano26b, Ano28i, Ano29c]. Important [Ano29c]. Improve [Rab27]. Including [Mor24]. Incomplete [Pea24d, RS21]. Indefinitely [Chu25, PJE29]. Independent [I.23]. Index [Ano24b, Mer28, Pea23b]. Indices [PT24]. Individual [Irw25a]. Individuals [Tip25]. Inequalities [Sho29]. Inference [NP28a, NP28b]. Infinite [Ney26b, Pea28a, Sop28]. Information [Ano20b, Ano22, Ano23c, Ano24d, Ano25d, Ano26d, Ano27e, Ano28j, Ano28k, Ano29d]. Inhabitants [Bux20].

Inheritance [Eld20, Eld23, War22]. Inherited [Sto23]. Inquiry [San24d]. Insulin [Rab27]. Integral [Wis25a, Wis25b]. Integrals [Cam24, Cam25b]. Intelligence [Eld23, P.23, PM27]. Inter [CGN23]. Inter-Relations [CGN23]. Interdependence [Pea27a]. Internal [HP29]. Interpolation [I.23, SE28]. Interpretation [NP28a, NP28b]. Interrelationships [MY23]. Interspecific [War24]. Interval [Mor21]. Investigation [Toc28]. Irwin [Pea25g]. Isserlis [P.23]. Issue [Ano24a].

J [Pea23a, Pea25g, Wis28a]. James [Pea25d]. Joint [Ano23b, PD21a, PD21b]. Judgments [Pea22a].

Katherine [Eld21]. King [Pea24j]. Kings [Tru28]. Kirkwall [Rei26]. Knee [Ano23b, PD21a, PD21b]. Knee-Joint [Ano23b, PD21a, PD21b]. Knowledge [HM26]. Korrelationsmessungen [And26, And27]. Korrelationstheorie [Ano26b]. Kr [Eld22].

L [P.23]. Laplace [Pea29c, Sim29]. Large [Chu25, PJE29, Wis25a, Wis25b]. Late [HM26]. Law [Irw27, Mer28]. Lebenden [Gum25]. Lebenserwartung [Gum25]. Left [HP29]. Lehrbuch [Ano25c]. Length [CGN23]. Lenses [Ush24]. Letter [Ano24a]. Life [Ano29c, GN25, Pea24e, Gum25]. Light [Pea25a]. Likeness [Pea26c]. Limiting [Nar23]. Limits [Pea28c]. Line [Pea21b]. Linear [Mus26b, Ney26a, Pea27b]. Living [Mah28, Pea24e, San24d, Til27]. Lock [Pea21a]. London [Mor26b]. Long [Ste29]. Lord [Pea28d].

Mabel [Eld21]. Macedonian [HM29]. Made [Pea25g]. Magnus [Rei26, Rei26]. Male [CGN23, San24d]. Males [GN25]. Manifold [Tap27]. Many [I.23]. Marked [Pea28c]. Martin [Ano25c]. Mary [Eld21, Pea28d]. Mathematical [Tch21, Wic25]. Mathematically [Nar23]. Mathematics [PM27]. Matter [Ano26a, Ano27b, Ano28b, Ano28a, Ano28c, Ano28d, Ano28f, Ano28e, Ano28g, Ano28h, Ano29a, Ano29b]. Mean [Kon29, Ney26b, Pea24h, Pea24e, Rid29]. Mean-Error [Pea24h]. Means [Chu26, Hal27b, HC28, Irw27, Yas25]. Measured [Woo29]. Measurement [HP29, Pea21c]. Measurements [Har26, HM29, Mah28, Mor27, MY23, Mus26a, Ore20, PS27, Til27, And26, And27]. median [Gum25]. Mediterranean [Bux20]. Members [Pea28c]. Memoir [Ano23b, P.25, Pea25g, Wic26, Woo28]. Men [HM29]. Mental [Art24, GTW25, Mus26a, SSK27]. Mesoblastic [Ush20]. Method [And26, And27, Pea21b, PE23, Pea28c, And23]. Methode [And23]. Methods [Mor26a, Pea25f]. Milk [Toc28]. Miss [P.25]. Mittleres [Gum25]. Mode [Yas26]. Modern [Wes25]. Moivre [Ano25b]. Moment [Pea28a, PJE29, Wis28b]. Moment-Coefficient [PJE29]. Moment-Coefficients [Pea28a]. Moments [Ano25a, Chu25, Fri25, Irw27, Pea24i, RS21, Rom25, San24c, Sho29, Tch21].

Moriori [Pea21d]. Mortality [GN25, Gum24]. Most [Ano29c]. Mr [Pea25g]. Mulberry [Ano28l]. Multiple [Cam25a, Hal27a, PE28, Pea25g, Tap27]. Multivariate [Wis28b]. Museum [Mor24]. Museums [Mor26b]. Mutually [Cam25a]. My [Pea25g].

Natives [Ore20]. Natural [Mor24, Pea23a]. Necessary [Pea23d]. Need [Mah28]. Negro [Til27]. Neolithic [Mor26b]. Nepalese [Mor24]. Neues [And23]. Newbold [P.25]. Non [Ney26a, PA29, Rid29, Woo29]. Non-Linear [Ney26a]. Non-Normal [PA29, Rid29]. Non-Significance [Woo29]. Normal [Ano21b, Ano25b, Mer28, Mil26, Pea24b, Pea26a, Pea27c, Pea28b, PA29, PJE29, Rid29, RS21, Tip25, Wis28b]. Note [Ano21a, Ano24c, Ano25a, Eld29, Hol24, Irw29, P.25, P.27, Pea21c, Pea23c, Pea24b, Pea24a, Pea24c, Pea24e, Pea24d, Pea25g, Pea25e, Pea26a, Pea27b, Pea27c, Pea29b, San24c, Ush20, Wis28a]. Notes [IR27, New25, Ney26a, Pea20b, Pea29a, Sto25]. Notice [P.23]. Notices [S.24]. Number [Pea28c].

O [Pea25g]. Observations [Irw25b]. Occurrences [Mor21]. One [Pea28a, Pea21c]. Only [Ano29c]. Order [PE28, Tho23]. Ordinate [Mil26]. Ordinates [Ano21b, Pea28a]. Oriental [Mor24]. Origin [Pea24b]. Origines [Sim29]. Orkney [Rei26]. Orthogonal [IR27]. Our [HM26]. Outlying [Irw25b].

Palaeopathology [S.24]. Paper [Ano24a, Ano24c, Irw29, Pea29b, Wis28a]. Parent [Cra29]. Parental [Smi22]. Part [Ano29c, Nar23, NP28a, NP28b, PD21b, Pea23a, Rho24]. Partial [Hal27a, New25, Tap27]. Patients [GTW25]. Pearson [Cra29, Irw27, Rom24]. Pedigree [Ush24]. Permille [Ano21b]. Personal [Pea22a]. Persons [Pea24e]. Physical [MY23, SSK27]. Pigmentation [Ush20]. Places [LP25]. Plana [Pea28b]. Point [Cam24, Fri25, Pea25e]. Political [Ano28i]. Polychoric [PP22]. Population [Chu25, Chu26, Cra29, Hal27b, HC28, Irw27, Ney26b, Pea24g, Pea26a, Pea27c, Pea28c, PJE29, Rho24, Sop28, SN25, Tip25, Wis28b]. Populations [PA28, PA29]. Portion [Mil26]. Portrait [Ano25b, Ano25c, Ano26c, Ano26b, Ano28i, Ano29c, P.24]. Portraits [Pea26d, Pea28d, Til23]. Position [Eld23]. Possible [Nar23]. Power [Pea25f]. Practical [Ano21a, Pea20a]. Pre [HP29]. Pre-Eminence [HP29]. Prediction [HP29]. Prehistoric [HM26, Mor25]. Preliminary [San24d]. Present [Ano24a, Eld23, HM26, Wis28a]. Previously [Mor27]. Price [Ano28i]. Print [Eld20]. Probability [Ano29c, Cam24, Cam25b, Wis25a]. Probable [Ano20a, Hal27b, Pea23b, Pea24a, Stu21, Yas26]. Problem [Ano21a, Irw25a, P.27, Pea20a, Pea25g, Rho24, Wic25, Wic26]. Problems [Ano29c, Pea27a, Wes25, E.25]. procedure [And23]. Product [PJE29, Wis28b]. Professor [Mor26a, Rom24]. Prove [Pea24e]. Psychical [Hol24]. Public [E.25]. Published [Irw29, Mor27]. Purposes [NP28a, NP28b].

Quadrature [Wis27]. Quantitative [Pea21c]. Queen [Pea28d]. Questions [Ano29c].

R [Eld21, Mor26a]. Race [PT24]. Racial [Pea26c, Ste29]. Random [Mor21]. Range [Pea28a, Pea26a, Tip25]. Ranking [Pea21c]. Ratio [Mil26, Mus26b, Pea23d, Pea27b, Rid29, Woo29]. Reaction [Har26]. Recently [Sto27]. Reconstruction [PS27, Ste29]. Records [Toc28]. Recurrence [Fri25]. Reference [Hoo26, Irw27, Mor26b, Ste29]. Regard [Ano24c, Eld23]. Regarded [Eld22]. Regression [Cam25a, Nar23, Ney26a, Pea21b, Ste29]. Reihenaus [And26, And27, And26, And27]. Reihenaus-Gleichungen [And26, And27]. Rejection [Irw25b]. Relation [Gra28, HP29, Hol24, P.23, Pea27c, SSK27]. Relations [CGN23]. Relationship [Art24, Har26, Pea24d, Pea25g, Pea26b]. Remains [Rei26]. Representation [Pea24f, Gum24]. Researches [Wis27]. Result [War22]. Review [E.25, Eld21, Eld22, I.23, Mor26a, P.23, R.23]. Richard [Ano28i]. Robert [Pea24j]. Rognvald [Rei26]. Roman [Mor25]. Romanovsky [Wis26]. Rome [Tru28]. Routine [Stu27]. Rudolf [Ano25c]. Ruml [Eld21].

S [Eld21, Mor26a]. Sa [Sim29]. Saint [Rei26]. same [Pea24g, Rho24]. Sample [CGN23, Pea28c]. Samples [Ano27c, Chu25, Chu26, Hal27b, HC28, Irw27, Ney26b, P.27, Pea24g, Pea25c, Pea26a, Pea27c, PA28, PA29, PJE29, Rho24, Rid29, Sop28, SN25, Tip25, Wis28b]. Sampling [Cra29, Pea25a, PM27, Pea29a, Pep29, Sun28]. School [P.23, Sto27]. Scotland [Mor26b, Pea24j]. Scots [Pea28d]. Scottish [Toc28]. Second [Pea21c, Wic26, Woo28]. Selection [Pea23a]. Selective [War22]. Semi [IR27]. Semi-Orthogonal [IR27]. Sensory [Art24]. Series [Ano25b, Ano25c, Ano26c, Ano26b, Ano28i, Ano29c, Cam25b, Mor24, Pea24f, Pea24i, Rom25]. Ses [Sim29]. Sesamoids [Ano23b, PD21a, PD21b]. Seven [LP25, Tru28]. Shakespeare [Ano23a]. Shaped [HC28]. Sight [Har26]. Significance [Woo29]. Silk [Ano28l]. Silk-Worm [Ano28l]. Similarities [Ano28l]. Simplification [Ano25a, San24a, San24b, San24c]. Sinistrality [WP27, Woo28]. Sir [P.24, Til23]. Size [Hal27b, Pea28c, SSK27]. Skew [Pea21b, PA29, Rho23, Rho25, Sop28, Wis27, Yas26]. Skull [HP29, Hoo26, Mor23, Pea21d, PD24, Pea24j, Pea26d, PS27, Pea28d, Til21, Til23, Til27, Mor26a]. Skulls [Mor26b, Mor27]. Small [Ano27c, Chu26, P.27, Pea25c, PA28, PA29, Rid29, Sop28, SN25]. Social [E.25]. Solid [RS21]. Solution [Pea27a]. Some [Pea29a, Rom24, Sto25, I.23]. Sound [Har26]. Space [Mor21]. Spasm [Sto23]. Spearman [Stu21]. Special [Hoo26, Irw27, Mor26b, Ste29]. Species [Ano28l, Pea27a]. Square [Kon29]. Squared [Chu25, Chu26]. Stabilitätsuntersuchungen [And26, And27]. Stability [PT24, And26, And27]. Standard [Chu25, Chu26, Kon29, Pea27c, Rid29, Smi22, Yas25]. Standard-Deviations [Chu25, Chu26]. Standardisation [Mah28]. Standpoint [Eld22]. State [HM26, Eld21]. Statement [Pea25g]. Statistical [NP28a, NP28b, Rab27, Sho29, Toc28]. Statistician [Ano28i]. Statistics [Ano21a, E.25, Pea20a, Wes25]. Stature [CGN23, Gra28, Ste29]. Stem [CGN23]. Sterbetafel [Gum24]. Stewart [Pea28d]. Street [Hoo26]. Strength [Art24]. Student [Ano24c]. Studies [Pep29]. Study [CGN23, Eld21, Hoo26, May23, Mor23, Mor24, Mor25, Mor26a, Mor26b, Mor27, Rab27, S.24, SSK27, Sto27, Til21, Wic25]. Subjected [Nar23]. Successive [Pea21b, Pea22a]. Such [Hal27b]. Sum [Pea24d]. Summary [Eld23]. Supplementary [Ano27d]. Supposed [Rho24]. Surface [Pea25b, Pea28b, Rho23, Rho25]. Surfaces [Cam25a, Pea24f]. survivors [Gum25]. Symmetrical [Mus26b, PA28, PA29, Wis25a]. System [Hal27a].

Table [Ano21b, LP25, Mil26, Gum24]. Tables [Ano27d, Woo29]. Taken [Pea26a, Tip25]. Takes [Hal27b]. Tamil [Mor26a]. Tasmanian [Mor27]. Tchouproff [Ano26b]. Terminals [Pea28a]. Terms [Pea21b, Pea24d]. Test [New25, NP28a, NP28b, Pea22b, Pea23c, Pea27c]. Testing [Ano24c]. Tests [Pea24g, Pea29a]. Tetrachoric [Ano27d, Hen22, LP25]. Their [E.25, Pea28d]. Theologian [Ano28i]. Theorem [Pea25d, Pea24c, Pea25a]. Theory [Ano27c, I.23, Irw25a, Pea23a, Pea25c, Pea27a, PM27, Pep29, SN25]. there [Pea28a]. Third [Hoo26]. Thomas [Til23, P.24, Wis27]. Thyroid [SSK27]. Tibetan [Mor23, Mor24]. Time [Har26, Mor21]. Times [HM26, Mor25, Mor26b]. Tolerance [Rab27]. Total [New25]. Tragedy [Pea28d]. Tuberculosis [Eld22]. Twenty [LP25]. Two [PE28, Pea24g, Pea29a, Rho24]. Type [Cra29, Irw27]. Types [Rom24].

U [HC28]. U-Shaped [HC28]. Universe [Tap27]. Universes [Rid29]. Unusual [PS23]. Use [Mor26a, NP28a, NP28b]. using [And23].

Values [Hal27b, Wis25a, Wis25b]. Variability [Eld29]. Variables [Dod27, Pea29a, I.23]. Variance [Ney26b]. Variate [And23, And26, And27, Hal27b, Hal27a, Pea21c, PE23]. Variate-Difference [And23]. Variates [Yas25]. Variation [Nar23]. Variations [Pea22a]. Varieties [Ano24c]. Various [Sho29]. Verfahren [And23]. Vital [CGN23, E.25, Hol24, MY23, Wes25]. Vol [Irw29]. Volume [Ano20b, Ano22, Ano23c, Ano24d, Ano25d, Ano26d, Ano27e, Ano28j, Ano28k, Ano29d, Wis28a]. Volumes [Ano24b].

Was [Pea21d]. Weight [CGN23, Eld29, Gra28]. Weights [GTW25]. Whether [Pea24g, Rho24]. Which [Nar23, Ano29c, Hal27b]. White [Pea21a]. White-Lock [Pea21a]. Wife [P.24]. Willis [Pea23a]. Wishart [Wis28a]. Within [PT24]. Women [Eld21, Mus26a]. Worm [Ano28l].

Writer [Ano28i].

XIX [Irw29]. XV [Ano24b]. XXa [Wis28a].

Year [GN25]. Years [Toc28]. Yield [Toc28]. York [Eld21].

Zoust [Ano23a].

References

Anderson:1923:NVA

[And23] Oskar Anderson. Über Ein Neues Verfahren Bei Anwendung Der “Variate-Difference”-Methode. (German) [On a new procedure using the “variate-difference” method]. Biometrika, 15(1/2):134–149, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331894.

Anderson:1926:ADV

[And26] Oskar Anderson. Über die Anwendung der Differenzenmethode (“Variate Difference Method”) bei Reihenaus-Gleichungen, Stabilitätsuntersuchungen und Korrelationsmessungen. (German) [On the application of the difference method (“Variate Difference Method”) with Reihenaus equations, stability analysis, and correlation measurements]. Biometrika, 18(3/4):293–320, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331955.

Anderson:1927:ADV

[And27] Oskar Anderson. Über die Anwendung der Differenzenmethode (“Variate Difference Method”) bei Reihenaus-Gleichungen, Stabilitätsuntersuchungen und Korrelationsmessungen. (German) [On the application of the difference method (“Variate Difference Method”) in Reihenaus equations, stability analysis, and correlation measurements]. Biometrika, 19(1/2):53–86, July 1927. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332177.

Anonymous:1920:PEF

[Ano20a] Anonymous. On the probable errors of frequency constants. Biometrika, 13(1):113–132, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331726.

Anonymous:1920:VI

[Ano20b] Anonymous. Volume information. Biometrika, 13(1):i–iv, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331719.

Anonymous:1921:NFP

[Ano21a] Anonymous. Note on the “Fundamental Problem of Practical Statistics”. Biometrika, 13(2/3):300–301, July 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331757.

Anonymous:1921:TON

[Ano21b] Anonymous. Table of ordinates of the normal curve for each permille of frequency. Biometrika, 13(4):426–428, October 1921. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331765.

Anonymous:1922:VI

[Ano22] Anonymous. Volume information. Biometrika, 14(1/2):i–iv, July 1922. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331854.

Anonymous:1923:ACZ

[Ano23a] Anonymous. Addenda: Concerning the “Zoust Shakespeare”. Biometrika, 15(1/2):76bis–76ter, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331890.

Anonymous:1923:AMS

[Ano23b] Anonymous. Addendum to a “Memoir on the Sesamoids of the Knee-Joint”. Biometrika, 14(3/4):409, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331820.

Anonymous:1923:VI

[Ano23c] Anonymous. Volume information. Biometrika, 15(1/2):i–iv, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331888.

Anonymous:1924:ELD

[Ano24a] Anonymous. Extract from a letter of Dr. E. I. Gumbel to the Editor of Biometrika concerning his paper in the present issue.

Biometrika, 16(3/4):411, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331717.

Anonymous:1924:IVX

[Ano24b] Anonymous. Index volumes I to XV. Biometrika, 16(3/4):1+3–39, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331718.

Anonymous:1924:NSR

[Ano24c] Anonymous. Note by “Student” with regard to his paper “On Testing Varieties of Cereals”. Biometrika, 16(3/4):411, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331716.

Anonymous:1924:VI

[Ano24d] Anonymous. Volume information. Biometrika, 16(1/2):i–vi, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331910.

Anonymous:1925:ENS

[Ano25a] Anonymous. Erratum: “Note on the Simplification of the Calculation of Abruptness Coefficients to Correct Crude Moments”. Biometrika, 17(1/2):ii, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332019. See [San24c].

Anonymous:1925:IBPa

[Ano25b] Anonymous. [illustration]: Biometrika portrait series: Abraham De Moivre: Discoverer of the “normal curve”. Biometrika, 17(1/2):??, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332020.

Anonymous:1925:IBPb

[Ano25c] Anonymous. [illustration]: Biometrika portrait series: Rudolf Martin, 1864–1925. “Lehrbuch der Anthropologie,” 1914. Biometrika, 17(3/4):??, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332076.

Anonymous:1925:VI

[Ano25d] Anonymous. Volume information. Biometrika, 17(1/2):i–iv, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332018.

Anonymous:1926:FM

[Ano26a] Anonymous. Front matter. Biometrika, 18(1/2):ii, July 1926. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332493.

Anonymous:1926:IBPb

[Ano26b] Anonymous. [illustration]: Biometrika portrait series: Alexander Tchouproff, 1847–1926. “Grundbegriffe und Grundprobleme der Korrelationstheorie,” 1925. Biometrika, 18(3/4):??, November 1926. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331951.

Anonymous:1926:IBPa

[Ano26c] Anonymous. [illustration]: Biometrika portrait series: Edmund Halley 1656–1742. Biometrika, 18(1/2):??, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332494.

Anonymous:1926:VI

[Ano26d] Anonymous. Volume information. Biometrika, 18(1/2):i–vii, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332492.

Anonymous:1927:ECF

[Ano27a] Anonymous. Errata et corrigenda. Biometrika, 19(1/2):vi, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332173. See [Chu26, Hoo26].

Anonymous:1927:FM

[Ano27b] Anonymous. Front matter. Biometrika, 19(1/2):ii, July 1927. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332172.

Anonymous:1927:FCT

[Ano27c] Anonymous. Further contributions to the theory of small samples. Biometrika, 19(3/4):441–442, December 1927. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331967.

Anonymous:1927:STD

[Ano27d] Anonymous. Supplementary tables for determining correlation from tetrachoric groupings. Biometrika, 19(3/4):354–404, December 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331964.

Anonymous:1927:VI

[Ano27e] Anonymous. Volume information. Biometrika, 19(1/2):i–v, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332171.

Anonymous:1928:BMb

[Ano28a] Anonymous. Back matter. Biometrika, 20B(1):??, July 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332243.

Anonymous:1928:BMa

[Ano28b] Anonymous. Back matter. Biometrika, 20A(1/2):??, July 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331946.

Anonymous:1928:BMc

[Ano28c] Anonymous. Back matter. Biometrika, 20B(2):??, September 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331950.

Anonymous:1928:BMd

[Ano28d] Anonymous. Back matter. Biometrika, 20A(3/4):??, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332123.

Anonymous:1928:FMb

[Ano28e] Anonymous. Front matter. Biometrika, 20B(1):??, July 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332241.

Anonymous:1928:FMa

[Ano28f] Anonymous. Front matter. Biometrika, 20A(1/2):??, July 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331937.

Anonymous:1928:FMc

[Ano28g] Anonymous. Front matter. Biometrika, 20B(2):??, September 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331948.

Anonymous:1928:FMd

[Ano28h] Anonymous. Front matter. Biometrika, 20A(3/4):??, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332109.

Anonymous:1928:IBP

[Ano28i] Anonymous. [illustration]: Biometrika portrait series: Dr Richard Price, 1723–1791 theologian, statistician, actuary and political writer. Biometrika, 20A(3/4):??, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332110.

Anonymous:1928:VIa

[Ano28j] Anonymous. Volume information. Biometrika, 20B(2):241–244, September 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331947.

Anonymous:1928:VIb

[Ano28k] Anonymous. Volume information. Biometrika, 20A(3/4):i–iv, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332108.

Anoochin:1928:SDS

[Ano28l] A. V. Anoochin. Similarities and differences in the species of the mulberry silk-worm. Biometrika, 20A(1/2):69–78, July 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331942.

Anonymous:1929:BM

[Ano29a] Anonymous. Back matter. Biometrika, 21(1/4):??, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332570.

Anonymous:1929:FM

[Ano29b] Anonymous. Front matter. Biometrika, 21(1/4):??, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332549.

Anonymous:1929:IBP

[Ano29c] Anonymous. [illustration]: Biometrika portrait series: “This Delicate Calculus applies to the Most Important Questions in Life, which are for the Most Part Only Problems in Probability”. Biometrika, 21(1/4):??, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332550.

Anonymous:1929:VI

[Ano29d] Anonymous. Volume information. Biometrika, 21(1/4):i–vi, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332548.

Arthur:1924:RBS

[Art24] William Arthur. On the relationship between strength of grip and certain mental and sensory characters. Biometrika, 16(3/4):299–327, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331711.

Burnside:1924:BF

[Bur24] W. Burnside. On Bayes’ formula. Biometrika, 16(1/2):189, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331919.

Buxton:1920:IEM

[Bux20] L. H. Dudley Buxton. The inhabitants of the Eastern Mediterranean. Biometrika, 13(1):92–112, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331725.

Camp:1924:PIP

[Cam24] Burton H. Camp. Probability integrals for the point binomial. Biometrika, 16(1/2):163–171, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331917.

Camp:1925:MCM

[Cam25a] Burton H. Camp. Mutually consistent multiple regression surfaces. Biometrika, 17(3/4):443–458, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332089.

Camp:1925:PIH

[Cam25b] Burton H. Camp. Probability integrals for a hypergeometric series. Biometrika, 17(1/2):61–67, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332024.

Cripps:1923:BSI

[CGN23] Lucy D. Cripps, Major Greenwood, and Ethel M. Newbold. A biometric study of the inter-relations of “vital capacity,” stature, stem length and weight in a sample of healthy male adults. Biometrika, 14(3/4):316–336, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331816.

Church:1925:MDS

[Chu25] A. E. R. Church. On the moments of the distribution of squared standard-deviations for samples of N drawn from an indefinitely large population. Biometrika, 17(1/2):79–83, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332026.

Church:1926:MSS

[Chu26] A. E. R. Church. On the means and squared standard-deviations of small samples from any population. Biometrika, 18(3/4):321–394, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331956.

See errata [Ano27a].

Craig:1929:SWP

[Cra29] Cecil Calvert Craig. Sampling when the parent population is of Pearson’s Type III. Biometrika, 21(1/4):287–293, December 1929. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332561.

Dodd:1927:CFC

[Dod27] Stuart C. Dodd. On criteria for factorising correlated variables. Biometrika, 19(1/2):45–52, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332176.

E:1925:BRB

[E.25] E. M. E. Book review: The Elements of Vital Statistics in Their Bearing on Social and Public Health Problems. Biometrika, 17(1/2):171–173, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332033.

Elderton:1920:IFP

[Eld20] Ethel M. Elderton. On the inheritance of the finger-print. Biometrika, 13(1):57–91, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331724.

Elderton:1921:BRB

[Eld21] Ethel M. Elderton. Book review: A Study of Women Delinquents in New York State by Mabel R. Fernald; Mary H. S. Hayes; Almena Dawley; Beardsley Ruml; Katherine Bement Davis. Biometrika, 13(2/3):305–308, July 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331759.

Elderton:1922:BRB

[Eld22] W. Palin Elderton. Book review: Is Tuberculosis to be Regarded from the Aetiological Standpoint as an Acute Disease of Childhood? by Kr. F. Andvord. Biometrika, 14(1/2):191–192, July 1922. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331861.

Elderton:1923:SPP

[Eld23] Ethel M. Elderton. A summary of the present position with regard to the inheritance of intelligence. Biometrika, 14(3/4):378–408, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331819.

Elderton:1929:NVG

[Eld29] Ethel M. Elderton. Note on variability in girls and boys (Glasgow) for height and weight. Biometrika, 21(1/4):429–430, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332568.

Frisch:1925:RFM

[Fri25] Ragnar Frisch. Recurrence formulae for the moments of the point binomial. Biometrika, 17(1/2):165–171, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332032.

Greenwood:1925:EMM

[GN25] Major Greenwood and E. M. Newbold. On the excess mortality of males in the first year of life. Biometrika, 17(3/4):327–342, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332084.

Gray:1928:RWS

[Gra28] Horace Gray. The relation of weight to stature, bi-cristal diameter and age. Biometrika, 20A(3/4):299–309, December 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332114.

Greenwood:1925:HWP

[GTW25] Rosa Greenwood, Cecily M. Thompson, and Hilda M. Woods. Heights and weights of patients in mental hospitals. Biometrika, 17(1/2):142–158, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332030.

Gumbel:1924:DSG

[Gum24] E. I. Gumbel. Eine Darstellung Der Sterbetafel. (German) [A representation of the mortality table]. Biometrika, 16(3/4):283–296, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331709.

Gumbel:1925:LMA

[Gum25] E. I. Gumbel. Lebenserwartung und Mittleres Alter der Lebenden. (German) [Life expectancy and median age of the survivors]. Biometrika, 17(1/2):173–175, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332034.

Hall:1927:MPC

[Hal27a] P. Hall. Multiple and partial correlation coefficients in the case of an n-fold variate system. Biometrika, 19(1/2):100–109, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332179.

Hall:1927:DMS

[Hal27b] Philip Hall. The distribution of means for samples of size N drawn from a population in which the variate takes values between 0 and 1, all such values being equally probable. Biometrika, 19(3/4):240–245, December 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331961.

Harmon:1926:DRB

[Har26] G. E. Harmon. On the degree of relationship between head measurements and reaction time to sight and sound. Biometrika, 18(1/2):207– 220, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332503.

Holzinger:1928:MSU

[HC28] K. J. Holzinger and A. E. R. Church. On the means of samples from a U-shaped population. Biometrika, 20A(3/4):361–388, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332118.

Henderson:1922:ETF

[Hen22] James Henderson. On expansions in tetrachoric functions. Biometrika, 14(1/2):157–185, July 1922. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331859.

Hooke:1926:PSO

[HM26] Beatrix G. E. Hooke and Geoffrey M. Morant. The present state of our knowledge of British craniology in late prehistoric and historic times. Biometrika, 18(1/2):99–104, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332497.

Hasluck:1929:MMM

[HM29] Margaret M. Hasluck and G. M. Morant. Measurements of Macedonian men. Biometrika, 21(1/4):322–336, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332564.

Holzinger:1924:NRV

[Hol24] Karl J. Holzinger. Note on the relation of vital capacity to certain psychical characters. Biometrika, 16(1/2):139–156, May 1924. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331915.

Hooke:1926:TSE

[Hoo26] Beatrix G. E. Hooke. A third study of the English skull with special reference to the Farringdon Street crania. Biometrika, 18(1/2):1– 55, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332495.

See errata [Ano27a].

Hoadley:1929:MID

[HP29] M. F. Hoadley and K. Pearson. On measurement of the internal diameters of the skull in relation: (i) to the prediction of its capacity, (II) to the “pre-eminence” of the left hemisphere. Biometrika, 21(1/4):85–123, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332553.

I:1923:BRB

[I.23] J. O. I. Book review: Some Formulas in the Theory of Interpolation of Many Independent Variables. Biometrika, 14(3/4):420, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331825.

Isserlis:1927:NCE

[IR27] L. Isserlis and V. Romanovsky. Notes on certain expansions in orthogonal and semi-orthogonal functions. Biometrika, 19(1/2):87–99, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332178.

Irwin:1925:FTF

[Irw25a] J. O. Irwin. The further theory of Francis Galton’s individual difference problem. Biometrika, 17(1/2):100–128, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332028.

Irwin:1925:CRO

[Irw25b] J. O. Irwin. On a criterion for the rejection of outlying observations. Biometrika, 17(3/4):238–250, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332079.

Irwin:1927:FDM

[Irw27] J. O. Irwin. On the frequency distribution of the means of samples from a population having any law of frequency with finite moments, with special reference to Pearson’s Type II. Biometrika, 19(3/4):225–239, December 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331960.

Irwin:1929:NPP

[Irw29] J. O. Irwin. Note on a paper published in Biometrika, vol. XIX. Biometrika, 21(1/4):431–432, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332569.

Kondo:1929:SEM

[Kon29] Tsutomu Kondo. On the standard error of the mean square contingency. Biometrika, 21(1/4):376–428, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332567.

Lee:1925:TFT

[LP25] Alice Lee and Karl Pearson. Table of the first twenty tetrachoric functions to seven decimal places. Biometrika, 17(3/4):343–354, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332085.

Mahalanobis:1928:NSM

[Mah28] P. C. Mahalanobis. On the need for standardisation in measurements on the living. Biometrika, 20A(1/2):1–31, July 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331938.

Maynard:1923:SHF

[May23] G. D. Maynard. A study in human fertility. Biometrika, 14(3/4):337– 354, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331817.

Merrill:1928:FDI

[Mer28] A. S. Merrill. Frequency distribution of an index when both the components follow the normal law. Biometrika, 20A(1/2):53–63, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331940.

Mills:1926:TRA

[Mil26] John P. Mills. Table of the ratio: Area to bounding ordinate, for any portion of normal curve. Biometrika, 18(3/4):395–400, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331957.

Morant:1921:ROS

[Mor21] G. Morant. On random occurrences in space and time, when followed by a closed interval. Biometrika, 13(4):309–337, October 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331760.

Morant:1923:FST

[Mor23] G. M. Morant. A first study of the Tibetan skull. Biometrika, 14(3/4):193–260, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331812.

Morant:1924:SCO

[Mor24] G. M. Morant. A study of certain oriental series of crania including the Nepalese and Tibetan series in the British Museum (natural history). Biometrika, 16(1/2):1–105, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331912.

Morant:1925:SEC

[Mor25] G. M. Morant. A study of Egyptian craniology from prehistoric to Roman times. Biometrika, 17(1/2):1–52, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332021.

Morant:1926:BRB

[Mor26a] G. M. Morant. Book review: The Use of Biometric Methods Applied to Craniology, Being a Critique of Professor Gordon Harrower’s “A Study of the Hokien and the Tamil Skull, by R. S. Edinburgh.”. Biometrika, 18(3/4):414–417, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331959.

Morant:1926:FSC

[Mor26b] G. M. Morant. A first study of the craniology of England and Scotland from Neolithic to early historic times, with special reference to the Anglosaxon skulls in London museums. Biometrika, 18(1/2):56– 98, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332496.

Morant:1927:SAT

[Mor27] G. M. Morant. A study of the Australian and Tasmanian skulls, based on previously published measurements. Biometrika, 19(3/4):417–440, December 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331966.

Morant:1929:CBC

[Mor29] G. M. Morant. A contribution to Basque craniometry. Biometrika, 21(1/4):67–84, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332552.

Musselman:1926:CHM

[Mus26a] J. R. Musselman. On the correlation of head measurements and mental agility: Women. Biometrika, 18(1/2):195–206, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332502.

Musselman:1926:LCR

[Mus26b] J. R. Musselman. On the linear correlation ratio in the case of certain symmetrical frequency distributions. Biometrika, 18(1/2):228–231, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332505.

Mumford:1923:IPM

[MY23] Alfred A. Mumford and Matthew Young. The interrelationships of the physical measurements and the vital capacity. Biometrika, 15(1/2):109–133, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331893.

Narumi:1923:GFB

[Nar23] Seimatsu Narumi. On the general forms of bivariate frequency distributions which are mathematically possible when regression and variation are subjected to limiting conditions: Part I. Biometrika, 15(1/2):77–88, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331891.

Newbold:1925:NET

[New25] Ethel M. Newbold. Notes on an experimental test of errors in partial correlation coefficients, derived from fourfold and biserial total coefficients. Biometrika, 17(3/4):251–265, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332080.

Neyman:1926:FNN

[Ney26a] J. Neyman. Further notes on non-linear regression. Biometrika, 18(3/4):257–262, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331953.

Neyman:1926:CMV

[Ney26b] J. Neyman. On the correlation of the mean and the variance in samples drawn from an “infinite” population. Biometrika, 18 (3/4):401–413, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331958.

Neyman:1928:UICa

[NP28a] J. Neyman and E. S. Pearson. On the use and interpretation of certain test criteria for purposes of statistical inference: Part I. Biometrika, 20A(1/2):175–240, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331945.

Neyman:1928:UICb

[NP28b] J. Neyman and E. S. Pearson. On the use and interpretation of certain test criteria for purposes of statistical inference: Part II. Biometrika, 20A(3/4):263–294, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332112.

Orensteen:1920:CCM

[Ore20] Myer M. Orensteen. Correlation of cephalic measurements in Egyptian born natives. Biometrika, 13(1):17–24, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331721.

P:1923:BRB

[P.23] K. P. Book review: Critical Notice: On the Relation Between Home Conditions and the Intelligence of School Children by L. Isserlis. Biometrika, 15(1/2):161–172, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331896.

P:1924:PST

[P.24] K. P. Portrait of Sir Thomas Browne and his wife. Biometrika, 16(1/2):204, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331926.

P:1925:NMN

[P.25] K. P. Note on miss Newbold’s memoir. Biometrika, 17(3/4):266–267, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332081.

P:1927:AHN

[P.27] K. P. Another “Historical Note on the Problem of Small Samples”. Biometrika, 19(1/2):207–210, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332184.

Pearson:1928:DFC

[PA28] Egon S. Pearson and N. K. Adyanthāya. The distribution of frequency constants in small samples from symmetrical populations. Biometrika, 20A(3/4):356–360, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332117.

Pearson:1929:DFC

[PA29] Egon S. Pearson and N. K. Adyanthāya. The distribution of frequency constants in small samples from non-normal symmetrical and skew populations. Biometrika, 21(1/4):259–286, December 1929. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332560.

Pearson:1921:SKJa

[PD21a] Karl Pearson and Adelaide G. Davin. On the sesamoids of the kneejoint. Biometrika, 13(2/3):133–175, July 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331752.

Pearson:1921:SKJb

[PD21b] Karl Pearson and Adelaide G. Davin. On the sesamoids of the kneejoint: Part II. Evolution of the sesamoids. Biometrika, 13(4):350–400, October 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331763.

Pearson:1924:BCH

[PD24] Karl Pearson and Adelaide G. Davin. On the biometric constants of the human skull. Biometrika, 16(3/4):328–363, December 1924.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331712.

Pearson:1923:VDM

[PE23] Karl Pearson and Ethel M. Elderton. On the variate difference method. Biometrika, 14(3/4):281–310, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331814.

P:1928:ACT

[PE28] K. P. and E. M. E. Appendix: On the correlation of two first order multiple correlation coefficients. Biometrika, 20A(3/4):310–313, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332115.

Pearson:1920:FPP

[Pea20a] Karl Pearson. The fundamental problem of practical statistics. Biometrika, 13(1):1–16, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331720.

Pearson:1920:NHC

[Pea20b] Karl Pearson. Notes on the history of correlation. Biometrika, 13(1):25–45, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331722.

Pearson:1921:NWL

[Pea21a] Karl Pearson. A new white-lock family (Barton’s case). Biometrika, 13(4):347–349, October 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331762.

Pearson:1921:GMD

[Pea21b] Karl Pearson. On a general method of determining the successive terms in a skew regression line. Biometrika, 13(2/3):296–300, July 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331756.

Pearson:1921:SNC

[Pea21c] Karl Pearson. Second note on the coefficient of correlation as determined from the quantitative measurement of one variate and the ranking of a second variate. Biometrika, 13(2/3):302–305, July 1921.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331758.

Pearson:1921:WSM

[Pea21d] Karl Pearson. Was the skull of the moriori artificially deformed? Biometrika, 13(4):338–346, October 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331761.

Pearson:1922:VPE

[Pea22a] Egon S. Pearson. On the variations in personal equation and the correlation of successive judgments. Biometrika, 14(1/2):23–102, July 1922. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331856.

Pearson:1922:TGF

[Pea22b] Karl Pearson. On the χ^{2} test of goodness of fit. Biometrika, 14(1/2):186–191, July 1922. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331860.

Pearson:1923:NSA

[Pea23a] Egon S. Pearson. Natural selection and the age and area theory of Dr J. C. Willis: Part I. Biometrika, 15(1/2):89–108, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331892.

Pearson:1923:PEC

[Pea23b] Egon S. Pearson. The probable error of a class-index correlation. Biometrika, 14(3/4):261–280, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331813.

Pearson:1923:FNT

[Pea23c] Karl Pearson. Further note on the χ^{2} test of goodness of fit. Biometrika, 14(3/4):418, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331823.

Pearson:1923:CNC

[Pea23d] Karl Pearson. On the correction necessary for the correlation ratio η. Biometrika, 14(3/4):412–417, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331822.

Pearson:1924:NAP

[Pea24a] Egon S. Pearson. Note on the approximation to the probable error of a coefficient of correlation. Biometrika, 16(1/2):196–198, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331922.

Pearson:1924:HNO

[Pea24b] Karl Pearson. Historical note on the origin of the normal curve of errors. Biometrika, 16(3/4):402–404, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331714.

Pearson:1924:NBT

[Pea24c] Karl Pearson. Note on Bayes’ Theorem. Biometrika, 16(1/2):190– 193, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331920.

Pearson:1924:NRI

[Pea24d] Karl Pearson. Note on the relationship of the incomplete Bfunction to the sum of the first p terms of the binomial (a + b)^{n}. Biometrika, 16(1/2):202–203, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331925.

Pearson:1924:NPE

[Pea24e] Karl Pearson. Note: To prove that the expectation of life at any age is greater than the mean age of all persons living greater than this age. Biometrika, 16(3/4):297–298, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331710.

Pearson:1924:CDH

[Pea24f] Karl Pearson. On a certain double hypergeometrical series and its representation by continuous frequency surfaces. Biometrika, 16(1/2):172–188, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331918.

Pearson:1924:DDT

[Pea24g] Karl Pearson. On the difference and the doublet tests for ascertaining whether two samples have been drawn from the same population. Biometrika, 16(3/4):249–252, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331706.

Pearson:1924:MEF

[Pea24h] Karl Pearson. On the mean-error of frequency distributions. Biometrika, 16(1/2):198–200, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331923.

Pearson:1924:MHS

[Pea24i] Karl Pearson. On the moments of the hypergeometrical series. Biometrika, 16(1/2):157–162, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331916.

Pearson:1924:SRB

[Pea24j] Karl Pearson. The skull of Robert The Bruce, King of Scotland, 1274–1329. Biometrika, 16(3/4):253–272, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331707.

Pearson:1925:BTE

[Pea25a] Egon S. Pearson. Bayes’ Theorem, examined in the light of experimental sampling. Biometrika, 17(3/4):388–442, December 1925. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332088.

Pearson:1925:FCB

[Pea25b] Karl Pearson. The fifteen constant bivariate frequency surface. Biometrika, 17(3/4):268–313, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332082.

Pearson:1925:FCT

[Pea25c] Karl Pearson. Further contributions to the theory of small samples. Biometrika, 17(1/2):176–200, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332035.

Pearson:1925:JBT

[Pea25d] Karl Pearson. James Bernoulli’s theorem. Biometrika, 17(3/4): 201–210, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332077.

Pearson:1925:NDA

[Pea25e] Karl Pearson. Note–the definition of the alveolar point. Biometrika, 17(1/2):53–56, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332022.

Pearson:1925:FPM

[Pea25f] Karl Pearson. On first power methods of finding correlation. Biometrika, 17(3/4):459–469, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332090.

Pearson:1925:MCB

[Pea25g] Karl Pearson. On the multiple correlation of brothers, being a note on Mr J. O. Irwin’s memoir, and on my statement of the application of Galton’s difference problem to the determination of the degree of relationship of brothers, made in August 1902. Biometrika, 17(1/2):129– 141, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332029.

Pearson:1926:FND

[Pea26a] Egon S. Pearson. A further note on the distribution of range in samples taken from a normal population. Biometrika, 18(1/2):173– 194, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332501.

Pearson:1926:AR

[Pea26b] Karl Pearson. On the avuncular relationship. Biometrika, 18 (1/2):231–232, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332506.

Pearson:1926:CRL

[Pea26c] Karl Pearson. On the coefficient of racial likeness. Biometrika, 18(1/2):105–117, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332498.

Pearson:1926:SPG

[Pea26d] Karl Pearson. On the skull and portraits of George Buchanan. Biometrika, 18(3/4):233–256, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331952.

Pearson:1927:ATD

[Pea27a] Egon S. Pearson. The application of the theory of differential equations to the solution of problems connected with the interdependence of species. Biometrika, 19(1/2):216–222, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332187.

Pearson:1927:FNL

[Pea27b] Egon S. Pearson. Further note on the “Linear Correlation Ratio”. Biometrika, 19(1/2):223–224, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332188.

Pearson:1927:NRG

[Pea27c] Karl Pearson. Note on the relation of the (P, χ^{2}) goodness of fit test to the distribution of standard deviations in samples from a normal population. Biometrika, 19(1/2):215, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332186.

Pearse:1928:CMC

[Pea28a] Gertrude E. Pearse. On corrections for the moment-coefficients of frequency distributions when there are infinite ordinates at one or both of the terminals of the range. Biometrika, 20A(3/4):314–355, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332116.

Pearson:1928:CGP

[Pea28b] Karl Pearson. The contribution of Giovanni Plana to the normal bivariate frequency surface. Biometrika, 20A(3/4):295–298, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332113.

Pearson:1928:MAL

[Pea28c] Karl Pearson. On a method of ascertaining limits to the actual number of marked members in a population of given size from a sample.

Biometrika, 20A(1/2):149–174, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331944.

Pearson:1928:SPH

[Pea28d] Karl Pearson. The skull and portraits of Henry Stewart, Lord Darnley, and their bearing on the tragedy of Mary, Queen of Scots. Biometrika, 20B(1):1–104, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332242.

Pearson:1929:SNS

[Pea29a] E. S. Pearson. Some notes on sampling tests with two variables. Biometrika, 21(1/4):337–360, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332565.

Pearson:1929:NDC

[Pea29b] Egon S. Pearson. Note on Dr Craig’s paper. Biometrika, 21 (1/4):294–302, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332562.

Pearson:1929:L

[Pea29c] Karl Pearson. Laplace. Biometrika, 21(1/4):202–216, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332557.

Pepper:1929:STS

[Pep29] Joseph Pepper. Studies in the theory of sampling. Biometrika, 21(1/4):231–258, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332559.

Pearson:1929:DFP

[PJE29] Karl Pearson, G. B. Jeffery, and Ethel M. Elderton. On the distribution of the first product moment-coefficient, in samples drawn from an indefinitely large normal population. Biometrika, 21(1/4):164–201, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332556.

Pearson:1927:MIS

[PM27] Karl Pearson and Margaret Moul. The mathematics of intelligence. The sampling errors in the theory of a generalised factor. Biometrika, 19(3/4):246–291, December 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331962.

Pearson:1922:PCC

[PP22] Karl Pearson and Egon S. Pearson. On polychoric coefficients of correlation. Biometrika, 14(1/2):127–156, July 1922. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331858.

Pearson:1923:UCD

[PS23] Karl Pearson and P. Stocks. On an unusual case of digital anomaly. Biometrika, 14(3/4):410–411, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331821.

Pearson:1927:FFR

[PS27] K. Pearson and Brenda N. Stoessiger. On further formulae for the reconstruction of cranial capacity from external measurements of the skull. Biometrika, 19(1/2):211–214, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332185.

Pearson:1924:SCI

[PT24] Karl Pearson and L. H. C. Tippett. On stability of the cephalic indices within the race. Biometrika, 16(1/2):118–138, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331914.

Pearson:1928:BC

[PT28] Karl Pearson and Charles F. Trustam. Biometry and chronology. Biometrika, 20A(3/4):241–262, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332111.

Pearson:1929:AD

[PU29] Karl Pearson and C. H. Usher. Albinism in dogs. Biometrika, 21 (1/4):144–163, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332555.

R:1923:BRB

[R.23] E. C. R. Book review: Frequency Arrays. Biometrika, 14(3/4):418– 419, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331824.

Rabinowitch:1927:DII

[Rab27] I. M. Rabinowitch. Does insulin improve carbohydrate tolerance in diabetes? (a statistical study). Biometrika, 19(3/4):405–416, December 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331965.

Reid:1926:RSM

[Rei26] R. W. Reid. Remains of Saint Magnus and Saint Rognvald, entombed in Saint Magnus Cathedral, Kirkwall, Orkney. Biometrika, 18(1/2):118–150, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332499.

Rhodes:1923:CSC

[Rho23] E. C. Rhodes. On a certain skew correlation surface. Biometrika, 14(3/4):355–377, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331818.

Rhodes:1924:PWT

[Rho24] E. C. Rhodes. On the problem whether two given samples can be supposed to have been drawn from the same population: Part I. Biometrika, 16(3/4):239–248, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331705.

Rhodes:1925:SCS

[Rho25] E. C. Rhodes. On a skew correlation surface. Biometrika, 17 (3/4):314–326, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332083.

Rider:1929:DRM

[Rid29] Paul R. Rider. On the distribution of the ratio of mean to standard deviation in small samples from non-normal universes. Biometrika, 21(1/4):124–143, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332554.

Romanovsky:1924:GST

[Rom24] V. Romanovsky. Generalisation of some types of the frequency curves of Professor Pearson. Biometrika, 16(1/2):106–117, May 1924. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331913.

Romanovsky:1925:MHS

[Rom25] V. Romanovsky. On the moments of the hypergeometrical series. Biometrika, 17(1/2):57–60, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332023.

Ritchie-Scott:1921:IMN

[RS21] A. Ritchie-Scott. The incomplete moments of a normal solid. Biometrika, 13(4):401–425, October 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331764.

S:1924:SPN

[S.24] P. S. The study of palaeopathology: Notices. Biometrika, 16(1/2):200–202, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331924.

Sandon:1924:ESCa

[San24a] Frank Sandon. Erratum: “Simplification of the Calculation of Abruptness Coefficients”. Biometrika, 16(1/2):ii, May 1924. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331911. See [San24c].

Sandon:1924:ESCb

[San24b] Frank Sandon. Erratum: “Simplification of the Calculation of Abruptness Coefficients”. Biometrika, 16(1/2):204, May 1924. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331927. See [San24c].

Sandon:1924:NSC

[San24c] Frank Sandon. Note on the simplification of the calculation of abruptness coefficients to correct crude moments. Biometrika, 16(1/2):193– 195, May 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331921.

See errata [San24a, San24b, Ano25a].

Sandon:1924:PID

[San24d] Frank Sandon. A preliminary inquiry into the density of the living male human body. Biometrika, 16(3/4):404–410, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331715.

Steffenson:1928:I

[SE28] J. F. Steffenson and W. P. E. Interpolation. Biometrika, 20A (3/4):425, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332122.

Shohat:1929:IMF

[Sho29] J. Shohat. Inequalities for moments of frequency functions and for various statistical constants. Biometrika, 21(1/4):361–375, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332566.

Simon:1929:OLS

[Sim29] L’Abbé G. A. Simon. Les origines de Laplace: Sa généalogie, ses Études. Biometrika, 21(1/4):217–230, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332558.

Smith:1922:SDF

[Smi22] Kirstine Smith. The standard deviations of fraternal and parental correlation coefficients. Biometrika, 14(1/2):1–22, July 1922. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331855.

Splawa-Neyman:1925:CTS

[SN25] J. Splawa-Neyman. Contributions to the theory of small samples drawn from a finite population. Biometrika, 17(3/4):472–479, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332092.

See discussion [Sop28].

Sophister:1928:DSS

[Sop28] Sophister. Discussion of small samples drawn from an infinite skew population. Biometrika, 20A(3/4):389–423, December 1928. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332119. See [SN25].

Stocks:1927:GAA

[SSK27] Percy Stocks, A. Vernon Stocks, and Mary N. Karn. Goitre in adolescence; an anthropometric study of the relation between the size of the thyroid gland and physical and mental development. Biometrika, 19(3/4):292–353, December 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331963.

Stevenson:1929:RDS

[Ste29] Paul Huston Stevenson. On racial differences in stature long bone regression formulae, with special reference to stature reconstruction formulae for the Chinese. Biometrika, 21(1/4):303–321, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332563.

Stocks:1923:FSI

[Sto23] Percy Stocks. Facial spasm inherited through four generations. Biometrika, 14(3/4):311–315, March 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331815.

Stocks:1924:CG

[Sto24] Percy Stocks. Cancer and goitre. Biometrika, 16(3/4):364–401, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331713.

Stocks:1925:SFN

[Sto25] Percy Stocks. Some further notes on cancer and goitre distributions. Biometrika, 17(1/2):159–165, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332031.

Stoessiger:1927:SBC

[Sto27] Brenda N. Stoessiger. A study of the Badarian crania recently excavated by the British School of Archaeology in Egypt. Biometrika, 19(1/2):110–150, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332180.

Student:1921:EDP

[Stu21] Student. An experimental determination of the probable error of Dr Spearman’s correlation coefficients. Biometrika, 13(2/3):263–282, July 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331754.

Student:1927:ERA

[Stu27] Student. Errors of routine analysis. Biometrika, 19(1/2):151–164, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332181.

Sun:1928:EFD

[Sun28] C. P. Sun. On the examination of final digits by experiments in artificial sampling. Biometrika, 20A(1/2):64–68, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331941.

Tappan:1927:PMC

[Tap27] M. Tappan. On partial multiple correlation coefficients in a universe of manifold characteristics. Biometrika, 19(1/2):39–44, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332175.

Tchouproff:1921:MEM

[Tch21] Al. A. Tchouproff. On the mathematical expectation of the moments of frequency distributions. Biometrika, 13(2/3):283–295, July 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331755.

Thomson:1923:HOA

[Tho23] Godfrey H. Thomson. On hierarchical order among correlation coefficients. Biometrika, 15(1/2):150–160, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331895.

Tildesley:1921:FSB

[Til21] M. L. Tildesley. A first study of the Burmese skull. Biometrika, 13(2/3):176–262, July 1921. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331753.

Tildesley:1923:STB

[Til23] M. L. Tildesley. Sir Thomas Browne: His skull, portraits, and ancestry. Biometrika, 15(1/2):1–76, August 1923. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331889.

Tildesley:1927:DCC

[Til27] Miriam L. Tildesley. Determination of the cranial capacity of the Negro from measurements on the skull or the living head. Biometrika, 19(1/2):200–206, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332183.

Tippett:1925:EIR

[Tip25] L. H. C. Tippett. On the extreme individuals and the range of samples taken from a normal population. Biometrika, 17(3/4):364–387, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332087.

Tocher:1928:IMY

[Toc28] J. F. Tocher. An investigation of the milk yield of dairy cows: Being a statistical analysis of the data of the Scottish milk records association for the years 1908, 1909, 1911, 1912, 1920 and 1923. Biometrika, 20B(2):105–240, September 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331949.

Trustam:1928:SKR

[Tru28] Charles F. Trustam. The seven Kings of Rome. Biometrika, 20A(3/4):424, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332121.

Usher:1920:HEA

[Ush20] C. H. Usher. Histological examination of an adult human Albino’s eyeball, with a note on mesoblastic pigmentation in foetal eyes. Biometrika, 13(1):46–56, October 1920. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331723.

Usher:1924:PCD

[Ush24] C. H. Usher. A pedigree of congenital dislocation of lenses. Biometrika, 16(3/4):273–282, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331708.

Warren:1922:IFR

[War22] Ernest Warren. Inheritance in the foxglove, and the result of selective breeding. Biometrika, 14(1/2):103–126, July 1922. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331857.

Warren:1924:IHD

[War24] Ernest Warren. On an interspecific hybrid of digitalis. Biometrika, 16(3/4):205–238, December 1924. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331704.

Westergaard:1925:MPV

[Wes25] Harald Westergaard. Modern problems in vital statistics. Biometrika, 17(3/4):355–363, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332086.

Wicksell:1925:CPM

[Wic25] S. D. Wicksell. The corpuscle problem: A mathematical study of a biometric problem. Biometrika, 17(1/2):84–99, June 1925. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332027.

Wicksell:1926:CPS

[Wic26] S. D. Wicksell. The corpuscle problem: Second memoir: Case of ellipsoidal corpuscles. Biometrika, 18(1/2):151–172, July 1926. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332500.

Wishart:1925:DLV

[Wis25a] John Wishart. Determination of ∫_{0}^{θ} cos^{n+1} θ dθ for large values of n, and its application to the probability integral of symmetrical frequency curves. Biometrika, 17(1/2):68–78, June 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332025.

Wishart:1925:FCI

[Wis25b] John Wishart. Further consideration of the integral ∫_{0}^{θ} cos^{n+1} θ dθ for large values of n. Biometrika, 17(3/4):469–472, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332091.

Wishart:1926:RGF

[Wis26] John Wishart. On Romanovsky’s generalised frequency curves. Biometrika, 18(1/2):221–228, July 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332504.

Wishart:1927:AQC

[Wis27] John Wishart. On the approximate quadrature of certain skew curves, with an account of the researches of Thomas Bayes. Biometrika, 19(1/2):1–38, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332174.

Wishart:1928:NPD

[Wis28a] J. Wishart. Note on the paper by Dr. J. Wishart in the present volume (XXa). Biometrika, 20A(3/4):424, December 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332120.

Wishart:1928:GPM

[Wis28b] John Wishart. The generalised product moment distribution in samples from a normal multivariate population. Biometrika, 20A(1/2): 32–52, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331939.

Woo:1928:DSH

[Woo28] T. L. Woo. Dextrality and sinistrality of hand and eye: Second memoir. Biometrika, 20A(1/2):79–148, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331943.

Woo:1929:TAS

[Woo29] T. L. Woo. Tables for ascertaining the significance or non-significance of association measured by the correlation ratio. Biometrika, 21 (1/4):1–66, December 1929. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332551.

Woo:1927:DSH

[WP27] T. L. Woo and Karl Pearson. Dextrality and sinistrality of hand and eye. Biometrika, 19(1/2):165–199, July 1927. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332182.

Yasukawa:1925:MSD

[Yas25] Kazutaro Yasukawa. On the means, standard deviations, correlations, and frequency distributions of functions of variates. Biometrika, 17(3/4):211–237, December 1925. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332078.

Yasukawa:1926:PEM

[Yas26] Kazutaro Yasukawa. On the probable error of the mode of skew frequency distributions. Biometrika, 18(3/4):263–292, November 1926. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331954.